# The evolution and detection of vector superradiant instabilities

### v1

@@ V.1 Gravitational wave signals
Although a Kerr BH does not radiate GW, the rotating vector condensate does. With only the dominant mode, the emitted GW is monochromatic, with a frequency equal to twice the energy of a vector particle in that mode. Nonetheless, there are other candidates, such as neutron stars, that could radiate monochromatic GW. Including the subdominant modes could break the degeneracy. Because the dominant and subdominant modes have a small difference in frequency, the coexistence of both modes results in a unique beat signature in the GW. Ref. Guo:2022mpr studied the beat signal from the BH-scalar-condensate. In this section, we study similar beat signals from the BH-vector-condensate systems.
@@
Since the GW signal is unavoidably mixed with the stochastic background, either large strength or more features are necessary to separate the signal. In both Figs. 10 and 11, strong GW requires both the BH initial BH mass and spin to be large. The rest of the parameter space can be explored with the help of the beat signal. Specifically, the strength of the beat is anti-correlated to the strength of the GW, which is expected with Eqs. (51) and (63). Therefore, the synergy of the GW strength and the beat feature could be important in searching for the BH-condensate systems in experiments.
In Fig. 10, one finds the $j=m=1$ modes radiate strong GW in a rather short time. It is proposed to observe the GW signal in the $j=1$ stage modes right after two BH merges Siemonsen:2022yyf . Comparably, the $j=2$ stage has a much longer GW emission time with the same parameters, longer than the $j=1$ stage by more than 7 orders of magnitude. The GW in the $j=2$ stage also has a much stronger beat signal, resulting from the larger mass ratio of the subdominant mode to the dominant one. Although the GW strength is weaker, the longer duration and stronger beat signal of the GW in the $j=2$ stage could facilitate the observation, providing a complementary method to search for ultralight vectors in the BH merger events.
For discussing detectability, the GW emission flux (30) should be converted to the GW strain amplitude. The details are available in Refs. Brito:2017zvb ; Brito:2017wnc ; Brito:2020lup ; Guo:2022mpr . We first ignore the frequency differences in Eq. (30). These quasi-monochromatic GWs have a frequency of $f_{\mathrm{s}}=\widetilde{\omega}/(2\pi)\simeq\mu/\pi$ in the source frame. The corresponding period is given by,
@@
where $r$ is the comoving distance. In the detector frame, the GW frequency is redshifted to $f_{\mathrm{d}}\equiv f_{\mathrm{s}}/(1+z)$. One could simply replace $f_{\text{s}}r$ in the formula by $f_{\text{d}}r_{\text{L}}$, where $r_{\text{L}}=r(1+z)$ is the luminosity distance. The $N_{\text{cycles}}$ represents the number of observed cycles, which is approximated by the minimum of the $f_{\mathrm{d}}T_{\mathrm{obs}}$ and $f_{\mathrm{s}}\tau_{\mathrm{GW}}$ in our calculation, where $T_{\mathrm{obs}}$ is the observation time. If the lifetime is short, the $N_{\text{cycles}}$ is determined by the source lifetime, otherwise it is constrained by the design of GW interferometers.
@@
Choosing an observation time $T_{\mathrm{obs}}=4\,\text{yr}$, a vector mass $10^{-16}\text{eV}$, and a redshift $z=1$, the number of observed cycles $N_{\text{cycles}}$ are approximately $10^{6}$.
When a dominant mode and a subdominant mode coexist, there is a characteristic beat signal due to the small frequency difference $\Delta\widetilde{\omega}$. The period of the beat in the source frame can be expressed as $T_{\mathrm{s,beat}}={2\pi}/{\Delta\widetilde{\omega}}$. Specifically, for the $j=m=1$ modes and $j=m=2$ modes, the periods are,
@@
In this case, the characteristic strain amplitude becomes Guo:2022mpr ,
@@
where the sum $\mathfrak{h}^{\mathrm{ave}}+\mathfrak{h}^{\mathrm{mod}}$ can be derived by changing every occurrence of $\left|U^{(\widetilde{\omega}_{i})}_{\widetilde{l}\widetilde{m}}\right|$ in Eq. (30) to $\left|U^{(\widetilde{\omega}_{i})}_{\widetilde{l}\widetilde{m}}\right|/\widetilde{\omega}_{i}^{2}$ with $i=1,2,3$. The $\mathfrak{h}^{\mathrm{ave}}$ is independent on time, while $\mathfrak{h}^{\mathrm{mod}}$ is time-modulated. If all sources have the same frequency, the modulated term $\mathfrak{h}^{\mathrm{mod}}$ is absent and Eq. (82) reduces to Eq. (76). Moreover, if the particle number of the dominant mode $N_{d}$ is much larger than that of the subdominant mode $N_{s}$, the $\mathfrak{h}^{\mathrm{ave}}$ and $\mathfrak{h}^{\mathrm{mod}}$ could be further simplified,
@@
We define the average characteristic strain as $\bar{h}_{\mathrm{c}}\equiv\sqrt{4N_{\text{cycles }}\mathfrak{h}^{\mathrm{ave}}/(5\mu^{2}r^{2})}$ while the beat characteristic strain as $\delta{h}_{\mathrm{c}}\equiv\max{(h_{\mathrm{c}})}-\min{(h_{\mathrm{c}})}$. In the limit of $N_{d}\gg N_{s}$, $\bar{h}_{c}$ reduces to Eq. (77) and,
@@
The fraction ${\left|U_{\widetilde{l}\widetilde{m}}^{(\widetilde{\omega}_{3})}\right|}/{\left|U_{\widetilde{l}\widetilde{m}}^{(\widetilde{\omega}_{1})}\right|}$ is $\sqrt{2}/4\approx 0.35$ for $j=1$ stage and $16/27\approx 0.59$ for $j=2$ stage. Thus, the beat modulation could reach $5\%$ even when $N_{s}/N_{d}$ is as small as $10^{-3}$.
@@
From Fig. 12, both the monochromatic GW and the GW beat in the $j=1$ stage have good detectability for DECIGO (from $10^{-16}$ to $10^{-13}$ eV), advanced LIGO (from $10^{-13}$ to $10^{-12}$ eV) and space-based detectors (from $10^{-19}$ to $10^{-15}$ eV). In the $j=2$ stage, DECIGO still has a very good potential while other detectors can only search for sources in our neighborhood. It is worth noting that the parameters in the above calculation are rather conservative. The detectability would be greatly improved with a larger initial BH mass and/or condensate mass.
@@ V.2 Black hole Regge trajectories
In addition to the direct detection of vector superradiance via gravitational waves, the distribution of BH spin and mass offers an indirect yet powerful method for detection. As previously discussed, a high-spin BH reduces its spin rapidly in the spin-down phase of each stage of evolution, eventually aligning with the Regge trajectory of the dominant modes, i.e. the (${}^{3}(j-1)_{j,j}^{0}$ modes. This alignment results in large forbidden regions in the BH ”Regge plot”—a plane that maps spin against mass. By scrutinizing the data from numerous BHs, we could identify both favored and unfavored vector mass ranges.
Figure 13: Evolution of BH spin-mass distribution with vector superradiance from solving Eqs. (41). The vector mass is fixed at $10^{-12}$ eV, and the initial mass of each mode is $10^{-5}$ of the initial BH mass. In total $10^{3}$ BHs are considered, distributed randomly in spin and log-uniformly in mass spanning from 3 $M_{\odot}$ to 100 $M_{\odot}$. The initial distribution is shown in panel (a). The distribution after 1 hour, 1 year, 10 years, $10^{6}$ years and $10^{10}$ years are shown in panels (b)-(f), respectively. In panels (b)-(f), the distribution of the previous moment is also shown in grey dots for comparison. In panels (b)-(f), the green curves from left to right correspond to the Regge trajectories of the dominant modes with $j=1,2,3,4,5$ from Eq. (38).
@@
After an hour of evolution, the BHs with high spins and initial mass $M_{0}\in(15,50)M_{\odot}$ descend to the ${}^{3}S_{1,1}^{0}$ Regge trajectory, as illustrated in Fig. 13(b). BHs with high spin but smaller mass require more time to align with the Regge trajectory. BHs positioned below this trajectory do not experience the $j=1$ stage, as the superradiant condition is not satisfied. After a year of evolution, shown in Fig. 13(c), BHs with high spin and large mass further descend to the ${}^{3}P_{2,2}^{0}$ trajectory, while BHs with high spin and smaller mass just have enough time to arrive at the ${}^{3}S_{1,1}^{0}$ trajectory. The later evolution follows the same pattern. With time proceeds, trajectories with larger $j$ consistently emerge on the higher mass end and those with smaller $j$ gradually become irrelevant, which is shown in Fig. 13(d)-(f).
@@
By comparing the observational data of BHs with the Regge plots obtained from different vector masses, we can constrain the mass of the vector field. Utilizing the 90 reported data from the LIGO-Virgo-KAGRA collaboration LIGOScientific:2018mvr ; LIGOScientific:2020ibl ; LIGOScientific:2021usb ; LIGOScientific:2021djp , with the corresponding parameter estimation samples from the online Gravitational-wave Transient Catalog  [^2] https://gwosc.org/eventapi/html/GWTC/, we calculate the median for BH masses and spins. The sources with median mass below $3M_{\odot}$ are excluded since they may potentially be neutron stars.
The resulting BH spin-mass distribution is plotted and compared with the Regge trajectories in Fig. 14. The trajectories are computed similarly to Fig. 13, but all initial BH spins are set as the limit value $a_{*\text{lim}}=0.998$ Thorne:1974ve . We present two typical vector masses $9\times 10^{-12}$ eV and $5\times 10^{-15}$ eV. For each mass, the darker color corresponds to $10^{6}$ years, while the lighter color represents $10^{10}$ years, approximately the age of the universe. This range is given by binary BH formation models LIGOScientific:2016vpg ; Arvanitaki:2016qwi , and we conservatively shift the lower bound from $10^{7}$ years to $10^{6}$ years.
@@
A strict treatment should employ the Bayesian hierarchical method, as demonstrated in the scalar case studies Ng:2019jsx ; Ng:2020ruv ; Fernandez:2019qbj ; Cheng:2022jsw , which is beyond the scope of this work.
@@ VI Summary
In Sec. V, we explore the direct and indirect detection approaches for vector superradiance, by studying the characteristic GW signals and by analyzing BH spin-mass distribution. For direct GW observations, we evaluate the characteristic strains of monochromatic waves for the ${}^{3}S_{1,1}^{0}$ mode and the ${}^{3}P_{2,2}^{0}$ mode, given in Eq. (78) and Eq. (79), respectively. The strain of the GW beat is in Eq. (85) in the limit where the mass ratio of the subdominant mode to the dominant mode is much less than 1, i.e., $N_{s}/N_{d}\ll 1$. Notably, the beat modulation can reach $5\%$ even when this ratio is as small as $10^{-3}$. Comparing with the characteristic noise strain of the current and projected GW detectors, for the $j=m=1$ case, both the monochromatic GWs and the GW beats have good detectability for DECIGO (from $10^{-16}$ to $10^{-13}$ eV), advanced LIGO (from $10^{-13}$ to $10^{-12}$ eV) and space-based detectors (from $10^{-19}$ to $10^{-15}$ eV). For the $j=m=2$ case, DECIGO still has a very good potential but other detectors can only search for the sources in our neighborhood. For the indirect detection, we first explain the process underlying the formation of forbidden regions in the Regge plot. By comparing the BH spin-mass data from the binary BH merger events to the forbidden regions with different vector masses, we found a conservative exclusion range for the vector mass: $5\times 10^{-15}\ \mathrm{eV}<\mu<9\times 10^{-12}\ \mathrm{eV}$.
@@ Acknowledgements
We give special thanks to Dr. G-R. Liang, Postdoctoral Fellow at China University of Mining and Technology, for his support and expertise during this research. S-S. Bao and Y-D. Guo and H. Zhang are supported by the National Natural Science Foundation of China (Grant No. 12075136) and the Natural Science Foundation of Shandong Province (Grant No. ZR2020MA094); N. Jia is supported by the National Natural Science Foundation of China (Grants Nos. 12147163 and 12175099); X. Zhang is supported by the National Natural Science Foundation of China (Grants Nos. 11975072 and 11835009), the National SKA Program of China (Grants Nos. 2022SKA0110200 and 2022SKA0110203), and the 111 Project (Grant No. B16009).

### v2

@@ V.1 Gravitational wave signals
Although a Kerr BH does not radiate GW, the rotating vector condensate does. With only the dominant mode, the emitted GW is monochromatic, with a frequency equal to twice the energy of a vector particle in that mode. Nonetheless, there are other candidates, such as neutron stars, that could radiate monochromatic GW. Including the subdominant modes could break the degeneracy. Because the dominant and subdominant modes have a small difference in frequency, the coexistence of both modes results in a unique beat signature in the GW. Ref. Guo:2022mpr studied the beat signal from the BH-scalar-condensate system. In this section, we study similar beat signals from the BH-vector-condensate systems.
@@
Since the GW signal is unavoidably mixed with the stochastic background, either large strength or more features are necessary to separate the signal. In both Figs. 10 and 11, a strong GW requires both the BH initial BH mass and spin to be large. The rest of the parameter space can be explored with the help of the beat signal. Specifically, the strength of the beat is anti-correlated to the strength of the GW, which is expected with Eqs. (51) and (63). Therefore, the synergy of the GW strength and the beat feature could be important in searching for the BH-condensate systems in experiments.
In Fig. 10, one finds the $j=m=1$ modes radiate strong GWs in a rather short time. It is proposed to observe the GW signal in the $j=1$ stage modes right after two BHs merge Siemonsen:2022yyf . Comparably, the $j=2$ stage has a much longer GW emission time with the same parameters, longer than the $j=1$ stage by more than 7 orders of magnitude. The GW in the $j=2$ stage also has a much stronger beat signal, resulting from the larger mass ratio of the subdominant mode to the dominant one. Although the GW strength is weaker, the longer duration and stronger beat signal of the GW in the $j=2$ stage could facilitate the observation, providing a complementary method to search for ultralight vectors in the BH merger events.
For discussing detectability, the GW emission flux (30) should be converted to the GW strain amplitude. The details are available in Refs. Brito:2017zvb ; Brito:2017wnc ; Brito:2020lup ; Guo:2022mpr . We first ignore the frequency differences in Eq. (30). These quasi-monochromatic GWs have a frequency of $f_{\mathrm{s}}=\widetilde{\omega}/(2\pi)\simeq\mu/\pi$ in the source frame. The corresponding period is given by
@@
where $r$ is the comoving distance. In the detector frame, the GW frequency is redshifted to $f_{\mathrm{d}}\equiv f_{\mathrm{s}}/(1+z)$. One could simply replace $f_{\text{s}}r$ in the formula by $f_{\text{d}}r_{\text{L}}$, where $r_{\text{L}}=r(1+z)$ is the luminosity distance. $N_{\text{cycles}}$ represents the number of observed cycles, which is approximated by the minimum of the $f_{\mathrm{d}}T_{\mathrm{obs}}$ and $f_{\mathrm{s}}\tau_{\mathrm{GW}}$ in our calculation, where $T_{\mathrm{obs}}$ is the observation time. If the lifetime is short, $N_{\text{cycles}}$ is determined by the source lifetime; otherwise, it is constrained by the design of GW interferometers.
@@
Choosing an observation time $T_{\mathrm{obs}}=4\,\text{yr}$, a vector mass $10^{-16}\text{eV}$, and a redshift $z=1$, the number of observed cycles $N_{\text{cycles}}$ is approximately $10^{6}$.
When a dominant mode and a subdominant mode coexist, there is a characteristic beat signal due to the small frequency difference $\Delta\widetilde{\omega}$. The period of the beat in the source frame can be expressed as $T_{\mathrm{s,beat}}={2\pi}/{\Delta\widetilde{\omega}}$. Specifically, for the $j=m=1$ modes and $j=m=2$ modes, the periods are
@@
In this case, the characteristic strain amplitude becomes Guo:2022mpr
@@
where the sum $\mathfrak{h}^{\mathrm{ave}}+\mathfrak{h}^{\mathrm{mod}}$ can be derived by changing every occurrence of $\left|U^{(\widetilde{\omega}_{i})}_{\widetilde{l}\widetilde{m}}\right|$ in Eq. (30) to $\left|U^{(\widetilde{\omega}_{i})}_{\widetilde{l}\widetilde{m}}\right|/\widetilde{\omega}_{i}^{2}$ with $i=1,2,3$. $\mathfrak{h}^{\mathrm{ave}}$ is independent of time, while $\mathfrak{h}^{\mathrm{mod}}$ is time-modulated. If all sources have the same frequency, the modulated term $\mathfrak{h}^{\mathrm{mod}}$ is absent and Eq. (82) reduces to Eq. (76). Moreover, if the particle number of the dominant mode $N_{d}$ is much larger than that of the subdominant mode $N_{s}$, $\mathfrak{h}^{\mathrm{ave}}$ and $\mathfrak{h}^{\mathrm{mod}}$ could be further simplified,
@@
We define the average characteristic strain as $\bar{h}_{\mathrm{c}}\equiv\sqrt{4N_{\text{cycles }}\mathfrak{h}^{\mathrm{ave}}/(5\mu^{2}r^{2})}$ and the beat characteristic strain as $\delta{h}_{\mathrm{c}}\equiv\max{(h_{\mathrm{c}})}-\min{(h_{\mathrm{c}})}$. In the limit of $N_{d}\gg N_{s}$, $\bar{h}_{c}$ reduces to Eq. (77) and
@@
The fraction ${\left|U_{\widetilde{l}\widetilde{m}}^{(\widetilde{\omega}_{3})}\right|}/{\left|U_{\widetilde{l}\widetilde{m}}^{(\widetilde{\omega}_{1})}\right|}$ is $\sqrt{2}/4\approx 0.35$ for the $j=1$ stage and $16/27\approx 0.59$ for the $j=2$ stage. Thus, the beat modulation could reach $5\%$ even when $N_{s}/N_{d}$ is as small as $10^{-3}$.
@@
From Fig. 12, both the monochromatic GW and the GW beat in the $j=1$ stage have good detectability for DECi-hertz Interferometer Gravitational wave Observatory (DECIGO) (from $10^{-16}$ to $10^{-13}$ eV), advanced LIGO (from $10^{-13}$ to $10^{-12}$ eV) and space-based detectors (from $10^{-19}$ to $10^{-15}$ eV). In the $j=2$ stage, DECIGO still has a very good potential while other detectors can only search for sources in our neighborhood. It is worth noting that the parameters in the above calculation are rather conservative. The detectability would be greatly improved with a larger initial BH mass and/or condensate mass.
@@ V.2 Black hole Regge trajectories
In addition to the direct detection of vector superradiance via gravitational waves, the distribution of BH spin and mass offers an indirect yet powerful method for detection. As previously discussed, a high-spin BH reduces its spin rapidly in the spin-down phase of each stage of evolution, eventually aligning with the Regge trajectory of the dominant modes, i.e., the (${}^{3}(j-1)_{j,j}^{0}$ modes. This alignment results in large forbidden regions in the BH ”Regge plot”—a plane that maps spin against mass. By scrutinizing the data from numerous BHs, we could identify both favored and unfavored vector mass ranges.
Figure 13: Evolution of BH spin-mass distribution with vector superradiance from solving Eqs. (41). The vector mass is fixed at $10^{-12}$ eV, and the initial mass of each mode is $10^{-5}$ of the initial BH mass. In total $10^{3}$ BHs are considered, distributed randomly in spin and log-uniformly in mass spanning from 3 $M_{\odot}$ to 100 $M_{\odot}$. The initial distribution is shown in panel (a). The distribution after 1 hour, 1 year, 10 years, $10^{6}$ years and $10^{10}$ years are shown in panels (b)-(f), respectively. In panels (b)-(f), the distribution of the previous moment is also shown in gray dots for comparison. In panels (b)-(f), the green curves from left to right correspond to the Regge trajectories of the dominant modes with $j=1,2,3,4,5$ from Eq. (38).
@@
After an hour of evolution, the BHs with high spins and initial mass $M_{0}\in(15,50)M_{\odot}$ descend to the ${}^{3}S_{1,1}^{0}$ Regge trajectory, as illustrated in Fig. 13(b). BHs with high spin but smaller mass require more time to align with the Regge trajectory. BHs positioned below this trajectory do not experience the $j=1$ stage, as the superradiant condition is not satisfied. After a year of evolution, shown in Fig. 13(c), BHs with high spin and large mass further descend to the ${}^{3}P_{2,2}^{0}$ trajectory, while BHs with high spin and smaller mass just have enough time to arrive at the ${}^{3}S_{1,1}^{0}$ trajectory. The later evolution follows the same pattern. As time proceeds, trajectories with larger $j$ consistently emerge on the higher mass end and those with smaller $j$ gradually become irrelevant, which is shown in Figs. 13(d)-(f).
@@
By comparing the observational data of BHs with the Regge plots obtained from different vector masses, we can constrain the mass of the vector field. Utilizing the 90 reported data from the LIGO-Virgo-KAGRA collaboration LIGOScientific:2018mvr ; LIGOScientific:2020ibl ; LIGOScientific:2021usb ; LIGOScientific:2021djp , with the corresponding parameter estimation samples from the online Gravitational-wave Transient Catalog, [^2] https://gwosc.org/eventapi/html/GWTC/ we calculate the median for BH masses and spins. The sources with median mass below $3M_{\odot}$ are excluded since they may potentially be neutron stars.
The resulting BH spin-mass distribution is plotted and compared with the Regge trajectories in Fig. 14. The trajectories are computed similarly to Fig. 13, but all initial BH spins are set as the limit value $a_{*\text{lim}}=0.998$ Thorne:1974ve . We present two typical vector masses $9\times 10^{-12}$ and $5\times 10^{-15}$ eV. For each mass, the darker color corresponds to $10^{6}$ years, while the lighter color represents $10^{10}$ years, approximately the age of the Universe. This range is given by binary BH formation models LIGOScientific:2016vpg ; Arvanitaki:2016qwi , and we conservatively shift the lower bound from $10^{7}$ years to $10^{6}$ years.
@@
In this analysis, we have ignored the large uncertainties of the BH spin data. A strict treatment should include the data uncertainty and employ the Bayesian hierarchical method, as demonstrated in the scalar case studies Ng:2019jsx ; Ng:2020ruv ; Fernandez:2019qbj ; Cheng:2022jsw , which is beyond the scope of this work.
@@ VI Summary
In Sec. V, we explore the direct and indirect detection approaches for vector superradiance, by studying the characteristic GW signals and by analyzing BH spin-mass distribution. For direct GW observations, we evaluate the characteristic strains of monochromatic waves for the ${}^{3}S_{1,1}^{0}$ mode and the ${}^{3}P_{2,2}^{0}$ mode, given in Eqs. (78) and (79), respectively. The strain of the GW beat is in Eq. (85) in the limit where the mass ratio of the subdominant mode to the dominant mode is much less than 1, i.e., $N_{s}/N_{d}\ll 1$. Notably, the beat modulation can reach $5\%$ even when this ratio is as small as $10^{-3}$. Comparing with the characteristic noise strain of the current and projected GW detectors, for the $j=m=1$ case, both the monochromatic GWs and the GW beats have good detectability for DECIGO (from $10^{-16}$ to $10^{-13}$ eV), advanced LIGO (from $10^{-13}$ to $10^{-12}$ eV) and space-based detectors (from $10^{-19}$ to $10^{-15}$ eV). For the $j=m=2$ case, DECIGO still has a very good potential but other detectors can only search for the sources in our neighborhood. For the indirect detection, we first explain the process underlying the formation of forbidden regions in the Regge plot. By comparing the BH spin-mass data from the binary BH merger events to the forbidden regions with different vector masses, we found a preliminary exclusion range for the vector mass: $5\times 10^{-15}\ \mathrm{eV}<\mu<9\times 10^{-12}\ \mathrm{eV}$.
@@ Acknowledgements
We give special thanks to Dr. Gui-Rong Liang, for making the collaboration possible and his continuous support during the research. S.-S. B. and Y.-D. G., and H. Z. are supported by the National Natural Science Foundation of China (Grant No. 12075136) and the Natural Science Foundation of Shandong Province (Grant No. ZR2020MA094); N. J. is supported by the National Natural Science Foundation of China (Grants No. 12147163 and No. 12175099); X. Z. is supported by the National Natural Science Foundation of China (Grant No. 12473001), the National SKA Program of China (Grants No. 2022SKA0110200 and No. 2022SKA0110203), and the 111 Project (Grant No. B16009).